40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 300 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  2001   Fri Sep 25 16:10:17 2009 JenneUpdateAdaptive FilteringSome progress on OAF, but more still to be done

[Jenne, Sanjit]

It seems now that we are able to get the OAF system to do a pretty good job of approximating the MC_L signal, but we can't get it to actually do any subtracting.  I think that we're not correctly setting the phase delay between the witness and the MC_L channels or something (I'm not sure though why we get a good filter match if the delay is set incorrectly, but we do get a good filter match for very different delay settings: 1, 5, 100, 1000 all seem to do equally well at adjusting the filter to match MC_L). 

The Matt Evans document in elog 395 suggests measuring the phase at the Nyquist frequency, and calculating the appropriate delay from that.  The sticking point with this is that we can't get test points for any channel which starts with C1:ASS.  I've emailed Alex to see what he can do about this.  Elog 1982 has a few words about how we're perhaps using a different awgtpman on the ass machine than we used to, which may be part of the problem. 

The golden plan, which in my head will work perfectly, is as follows: Alex will fix the testpoint problem, then Sanjit and I will be able to measure the phase between our OAF signal and the incoming MC_L signal, we will be able to match them as prescribed in the Matt Evans document, and then suddenly the Adaptive Filtering system will do some actual subtracting!

The plot below shows the Reference MC_L without any OAF system (black), the output of the OAF (green), and the 'reduced' MC_L (red).  As you can see, the green trace is doing a pretty good job of matching the black one, but the red trace isn't getting reduced at all.

Attachment 1: OAF_Running_25Sept2009.jpg
OAF_Running_25Sept2009.jpg
  2000   Thu Sep 24 21:04:15 2009 JenneUpdateMOPAIncreasing the power from the MOPA

[Jenne, Rana, Koji]

Since the MOPA has been having a bad few weeks (and got even more significantly worse in the last day or so), we opened up the MOPA box to increase the power.  This involved some adjusting of the NPRO, and some adjusting of the alignment between the NPRO and the Amplifier.  Afterward, the power out of the MOPA box was increased.  Hooray! 

Steps taken:

0.  Before we touched anything, the AMPMON was 2.26, PMC_Trans was 2.23, PSL-126MOPA_126MON was 152 (and when the photodiode was blocked, it's dark reading was 23).

1.  We took off the side panel of the MOPA box nearest the NPRO, to gain access to the potentiometers that control the NPRO settings.  We selectively changed some of the pots while watching PSL-126MOPA_126MON on Striptool.

2.  We adjusted the pot labeled "DTEMP" first. (You have to use a dental mirror to see the labels on the PCB, but they're there). We went 3.25 turns clockwise, and got the 126MON to 158. 

3. To give us some elbow room, we changed the PSL-126MOPA_126CURADJ from +10.000 to 0.000 so that we have some space to move around on the slider.  This changed 126MON to 142. The 126MOPA_CURMON was at 2.308.

4.  We tried adjusting the "USR_CUR" pot, which is labeled "POWER" on the back panel of the NPRO (you reach this pot through a hole in the back of the NPRO, not through the side which we took off, like all the other pots today).  This pot did nothing at all, so we left it in its original position.  This may have been disabled since we use the slider.

5.  We adjusted the CUR_SET pot, and got the 126MON up to 185.  This changed the 126MOPA_CURMON to 2. 772 and the AMPMON to 2.45

We decided that that was enough fiddling with the NPRO, and moved on to adjusting the alignment into the Amplifier.

6.  We teed off of the AMPMON photodiode so that we could see the DC values on a DMM.  When we used a T to connect both the DMM and the regular DAQ cable, the DMM read a value a factor of 2 smaller than when the DMM was connected directly to the PD.  This shouldn't happen.....it's something on the to-fix-someday list.

7.  Rana adjusted the 2 steering mirrors immediately in front of the amplifier, inside the MOPA box.  This changed the DMM reading from its original 0.204 to 0.210, and the AMPMON reading from 2.45 to 2.55. While this did help increase the power, the mirrors weren't really moved very much.

8.  We then noticed that the beam wasn't really well aligned onto the AMPMON PD.  When Rana leaned on the MOPA box, the PD's reading changed.  So we moved the PD a little bit to maximize its readings.  After this, the AMPMON read 2.68, and the DMM read 0.220.

9.  Then Rana adjusted the 2 waveplates in the path from the NPRO to the Amplifier.  The first waveplate in the path didn't really change anything.  Adjusting the 2nd waveplate gave us an AMPMON of 2.72, and a DMM reading of 0.222.

10.  We closed up the MOPA box, and locked the PMC.  Unfortunately, the PMC_Trans was only 1.78, down from the 2.26 when we began our activities.  Not so great, considering that in the end, the MOPA power went up from 2.26 to 2.72.

11.  Koji and I adjusted the steering mirrors in front of the PMC, but we could not get a transmission higher than 1.78.

12.  We came back to the control room, and changed the 126MOPA_126CURADJ slider to -2.263 which gives a 126MOPA_CURMON to 2.503.  This increased PMC_TRANS up to 2.1. 

13.  Koji did a bit more steering mirror adjustment, but didn't get any more improvement.

14.  Koji then did a scan of the FSS SLOW actuator, and found a better temperature place (~ -5.0)for the laser to sit in.  This place (presumably with less mode hopping) lets the PMC_TRANS get up to 2.3, almost 2.4.  We leave things at this place, with the 126MOPA_126CURADJ slider at -2.263. 

Now that the MOPA is putting out more power, we can adjust the waveplate before the PBS to determine how much power we dump, so that we have ~constant power all the time.

 

Also, the PMCR view on the Quad TVs in the Control Room has been changed so it actually is PMCR, not PMCT like it has been for a long time.

  1999   Thu Sep 24 20:17:05 2009 ranaSummaryLSCComparison of Material Properties for the new RFPD Mounts
  Steel Brass Aluminum Delrin
Density (kg/m^3) 7850 8500 2700 1420
CTE (ppm/C) 12 19 23 100

Young's

Modulus

(GPa)

200 110 69 2
Hardness        
Color grey gold light silver any

 

  1998   Thu Sep 24 19:35:20 2009 ranaHowToPhotos40m Google account

I've created a 40m Google account. Please post all the 40m related photos to this site. If you don't already have it, download Picasa to make this easier.

40m Installation Photos">

the password is in the usual password place.

  1997   Thu Sep 24 15:45:27 2009 robUpdateIOOMC OLG

I measured the mode cleaner open loop gain with the HP3563A

The UGF is 64kHz, phase margin is 28 deg.

  1996   Wed Sep 23 20:02:11 2009 JenneAoGComputersGremlins in the RFM

Quote:

A cosmic ray struck the RFM in the framebuilder this afternoon, causing hours of consternation.  The whole FE system is just now coming back up, and it appears the mode cleaner is not coming back to the same place (alignment).

 

rob, jenne

 Jenne, Rana, Koji

The mode cleaner has been realigned, using a combination of techniques.  First, we used ezcaservo to look at C1:SUS-MC(1,3)_SUS(DOF)_INMON and drive C1:SUS-MC(1,3)_(DOF)_COMM, to put the MC1 and MC3 mirrors back to their DriftMon values.  Then we looked at the MC_TRANS_SUM on dataviewer and adjusted the MC alignment sliders by hand to maximize the transmission.  Once the transmission was reasonably good, we saw that the spot was still a little high, and the WFS QPDs weren't centered.  So Koji and I went out and centered the WFS, and now the MC is back to where it used to be.  The MC_TRANS QPD looks nice and centered, so the pointing is back to where it used to be.

  1995   Wed Sep 23 19:36:41 2009 ranaUpdatePSLRC temperature performance

This first plot shows the RC temperature channels' performance from 40 days ago, before we disabled the MINCO PID controller. Although RCTEMP is supposed to be the out of loop sensor, what we really care about is the cavity length and so I've plotted the SLOW. To get the SLOW on the same scale, I've multiplied the channel by 10 and then adjusted the offset to get it on the same scale.

 The second plot shows a period after that where there is no temperature control of the can at all. Same gain scaling has been applied to SLOW as above, so that instead of the usual 1 GHz/V this plot shows it in 0.1 GHz/V.

The third plot shows it after the new PID was setup.

Summary: Even though the PID loop has more gain, the true limit to the daily fluctuations in the cavity temperature and the laser frequency are due to the in-loop sensors measuring the wrong thing. i.e. the out-of-loop temperature is too different from the in-loop sensor. This can possibly be cured with better foam and better placement of the temperature sensors. Its possible that we're now just limited by the temperature gradients on the can.

Attachment 1: Untitled.png
Untitled.png
Attachment 2: Untitled.png
Untitled.png
Attachment 3: e.png
e.png
  1994   Wed Sep 23 17:32:37 2009 robAoGComputersGremlins in the RFM

A cosmic ray struck the RFM in the framebuilder this afternoon, causing hours of consternation.  The whole FE system is just now coming back up, and it appears the mode cleaner is not coming back to the same place (alignment).

 

rob, jenne

  1993   Fri Sep 18 16:26:02 2009 steveSummaryPSLNeslab chiller is OK

Rob found puddles of water very close to the chiller during lunch time. We raised the unit and took the side cover off. All surfaces were dry and the water level in the tub normal.

Later on we discovered that one of the Vons distilled water bottle was leaking. Jenne and I checked for excess amount of condensing water droplets inside the MOPA box.

On the bare,not insulated tubing and valve are loaded with droplets of water. Relative humidity is 44% at 24 C and HEPA filter speed set to 80 V in the enclosure.

 

  1992   Fri Sep 18 16:05:08 2009 JenneOmnistructurePSLwater under the laser chiller

Quote:

rob, koji, steve

We noticed some water (about a cup) on the floor under the NESLAB chiller today.  We put the chiller up on blocks and took off the side panel for a cursory inspection, but found no obvious leaks.  We'll keep an eye on it.

 The culprit has been found:  One of the bottles of chiller water had a tiny leak in it, and apparently the floor is sloped just right to make it look like the water had been coming from under the chiller.  All is well again in the world of chilled water.

  1991   Fri Sep 18 14:25:00 2009 robOmnistructurePSLwater under the laser chiller

rob, koji, steve

We noticed some water (about a cup) on the floor under the NESLAB chiller today.  We put the chiller up on blocks and took off the side panel for a cursory inspection, but found no obvious leaks.  We'll keep an eye on it.

  1990   Thu Sep 17 15:05:47 2009 robUpdateComputersawgtpman on c1omc failing to start

Quote:

[root@c1omc controls]# /opt/gds/awgtpman -2 &
[1] 16618
[root@c1omc controls]# mmapped address is 0x55577000
32 kHz system
Spawn testpoint manager
no test point service registered
Test point manager startup failed; -1

[1]+  Exit 1                  /opt/gds/awgtpman -2

 

 

 

 

 

 

This turned out to be fallout from the /cvs/cds transition.  Remounting and restarting fixed it.

  1989   Thu Sep 17 14:17:04 2009 robUpdateComputersawgtpman on c1omc failing to start

[root@c1omc controls]# /opt/gds/awgtpman -2 &
[1] 16618
[root@c1omc controls]# mmapped address is 0x55577000
32 kHz system
Spawn testpoint manager
no test point service registered
Test point manager startup failed; -1

[1]+  Exit 1                  /opt/gds/awgtpman -2

 

 

 

  1988   Wed Sep 16 11:58:11 2009 JenneUpdateAdaptive FilteringNew Filters for Adaptive Filtering

When Sanjit and I were looking at the adaptive filtering system on Monday and Friday, we noticed that turning on the Accelerometers (which had been used in the past) seemed to do good things, but that turning on the seismometers (which I just put into the system last week) made the OAF output integrate up.  Rana pointed out that this is an indication of a missing high pass filter.  And indeed, when I put the seismometers in, I neglected to copy the high pass filter at low frequencies, and the low pass at 64Hz from the accelerometer path to the seismometer path.  The accelerometers had a HP at 1Hz, which is okay since they don't really do useful things down to the mHz level.  I gave all of the seismometers HP at 1mHz.  These are now in the filter banks in the ASS_TOP_PEM screen.  The accelerometers are on channels 15, 16, 17, 18, 19, 20 and the seismometers are on channels 2, 3, 4, 10, 11, 12, 24.

I now need to modify the upass script to turn these filters on before doing adaptive filtering.

  1987   Tue Sep 15 15:46:05 2009 steveUpdatePEMPEM and VAC

FSS_RMTEMP is moving up and  daily fluctuations  are  less . 120 and 16 days plots are below.

Attachment 1: 120dpem.jpg
120dpem.jpg
Attachment 2: 16dpem.jpg
16dpem.jpg
  1986   Sat Sep 12 15:40:15 2009 ranaUpdatePSLRC response v. can temperature

I stepped the RC can temperature to see the response in the laser frequency. This gives a true measure of the thermal time constant of the RC. Its ~4 hours.

Since the RCPID screen now has a setpoint field, I can remotely type in 1 deg steps. The NPRO SLOW actuator locks the NPRO to the RC at long time scales and so we can use C1:PSL-FSS_SLOWDC to measure the RC length. By knowing what the step response time constant is, we can estimate the transfer function from can temperature to frequency noise and thereby make a better heater circuit.

 Does the observed temperature shift make any sense? Well, John Miller and I measured the SLOW calibration to be 1054 +/- 30 MHz / V.

We know that the thermal expansion coefficient of fused silica, alpha = 5.5 x 10-7 (dL/L)/deg. So the frequency shift ought to be alpha * c / lambda = 155 MHz / deg.

Instead we see something like 110 MHz / deg. We have to take more data to see if the frequency shift will actually asymptote to the right value or not. If it doesn't, one possibility is that we are seeing the effect of temperature on the reflection phase of the mirror coatings through the dn/dT and the thermal expansion of the dielectric layers. I don't know what these parameters are for the Ta2O5 layers.

A more useful measure of the frequency noise can be gotten by just looking at the derivative in the first 30 minutes of the step, since that short time scale is much more relevant for us. Its 0.04 V / hour / (2 deg) =>  860 (Hz/s)/deg.

In the frequency domain this comes out to be dnu/dT = 860 Hz/deg @ 0.16 Hz or dnu/dT = 137 *(1/f) Hz / deg.

Our goal for the reference cavity frequency noise is 0.01 * (1/f) Hz/rHz. So the temperature noise of the can needs to be < 0.1 mdeg / rHz.

Attachment 1: Picture_2.png
Picture_2.png
Attachment 2: Untitled.pdf
Untitled.pdf
  1985   Fri Sep 11 17:11:15 2009 SanjitUpdateASSOAF: progress made

[Jenne & Sanjit]

Good news: We could successfully send filtered output to MC1 @ SUS.

We used 7 channels (different combinations of 3 seismometer and six accelerometer)

We tried some values of \mu (0.001-0.005) & gain on SUS_MC1_POSITION:MCL and C1ASS_TOP_SUS_MC1 (0.1-1).

C1:ASS-TOP_SUS_MC1_INMON is huge (soon goes up to few times 10000), so ~0.1 gains at two places bring it down to a reasonable value.

Bad news: no difference between reference and filtered IOO-MC_L power spectra so far.

Plan of action: figure out the right values of the parameters (\mu, \tau, different gains, and may be some delays), to make some improvement to the spectra.

 ** Rana: there's no reason to adjust any of the MCL gains. We are not supposed to be a part of the adaptive algorithm.

  1984   Fri Sep 11 17:07:45 2009 JenneUpdateAdaptive FilteringMinor changes to ASS_TOP_PEM screen.

There was some uncertainty as to which channels were being input into the Adaptive Filtering screen, so I checked it out to confirm.  As expected, the rows on the ASS_TOP_PEM screen directly correspond to the BNC inputs on the PEM_ADCU board in the 1Y6 (I think it's 6...) rack.  So C1:ASS-TOP_PEM_1_INMON corresponds to the first BNC (#1) on the ADCU, etc. 

After checking this out, I put text tags next to all the inputs on the ASS_TOP_PEM screen for all of the seismometers (which had not been there previously).  Now it's nice and easy to select which witness channels you want to use for the adaptation.

  1983   Thu Sep 10 18:25:15 2009 ranaUpdatePSLc1psl rebooted for new RCPID database settings

I added a new database record (C1:PSL-FSS_RCPID_SETPOINT) to allow for changing of the RC setpoint while the loop is on. This will enable us to step the can's temperature and see the result in the NRPO's SLOWDC.

 

  1982   Thu Sep 10 17:47:25 2009 JenneUpdateComputerschanges to the startass scripts

[Rana, Jenne]

While I was mostly able to restart the c1ass computer earlier today, the filter banks were acting totally weird.  They were showing input excitations when we weren't putting any, and they were showing that the outputs were all zero, even though the inputs were non-zero and the input and the output were both enabled. The solution to this ended up being to use the 2nd to last assfe.rtl backup file.  Rana made a symbolic link from assfe.rtl to the 2nd to last backup, so that the startup.cmd script does not need to be changed whenever we alter the front end code.

The startup_ass script, in /caltech/target/gds/ which, among other things, starts the awgtpman was changed to match the instructions on the wiki Computer Restart page.  We now start up the /opt/gds/awgtpman .  This may or may not be a good idea though, since we are currently not able to get channels on DTT and Dataviewer for the C1:ASS-TOP_PEM channels.  When we try to run the awgtpman that the script used to try to start ( /caltech/target/gds/bin/ ) we get a "Floating Exception". We should figure this out though, because the /opt/gds/awgtpman does not let us choose 2kHz as an option, which is the rate that the ASS_TOP stuff seems to run at.

The last fix made was to the screen snapshot buttons on the C1:ASS_TOP screen.  When the screen was made, the buttons were copied from one of the other ASS screens, so the snapshots saved on the ASS_TOP screen were of the ASS_PIT screen.  Not so helpful.  Now the update snapshot button will actually update the ASS_TOP snapshot, and we can view past ASS_TOP shots.

  1981   Thu Sep 10 15:55:44 2009 JenneUpdateComputersc1ass rebooted

c1ass had not been rebooted since before the filesystem change, so when I was sshed into c1ass I got an error saying that the NFS was stale.  Sanjit and I went out into the cleanroom and powercycled the computer.  It came back just fine.  We followed the instructions on the wiki, restarting the front end code, the tpman, and did a burt restore of c1assepics. 

  1980   Wed Sep 9 10:13:31 2009 steveBureaucracyGeneralthe use of FAX machine, scanner & toaster

 

 You said that the use of FAXST was forbidden for phds and graduate students. I had to swear on the promise of not ever buying an other FAXST

  1979   Tue Sep 8 20:25:03 2009 JenneOmnistructureDMFDMF restarted

Quote:

I (think I) restarted DMF.  It's on Mafalda, running in matlab (not the complied version which Rana was having trouble with back in the day).  To start Matlab, I did "nohup matlab", ran mdv_config, then started seisBLRMS.m running.  Since I used nohup, I then closed the terminal window, and am crossing my fingers in hopes that it continues to work.  I would have used Screen, but that doesn't seem to work on Mafalda.

 Just kidding. That plan didn't work.  The new plan: I started a terminal window on Op540, which is ssh-ed into Mafalda, and started up matlab to run seisBLRMS.  That window is still open. 

Because Unix was being finicky, I had to open an xterm window (xterm -bg green -fg black), and then ssh to mafalda and run matlab there.  The symptoms which led to this were that even though in a regular terminal window on Op540, ssh-ed to mafalda, I could access tconvert, I could not make gps.m work in matlab.  When Rana ssh-ed from Allegra to Op540 to Mafalda and ran matlab, he could get gps.m to work.  So it seems like it was  a Unix terminal crazy thing. Anyhow, starting an xterm window on Op540m and ssh-ing to mafalda from there seemed to work.

Hopefully this having a terminal window open and running DMF will be a temporary solution, and we can get the compiled version to work again soon.

  1978   Tue Sep 8 20:15:33 2009 rana, jenneSummaryPSLRC temperature servo: Heater Voltage noise

We measured the voltage noise of the heater used to control the RC can temperature. It is large.

TEK00074.PNG

The above scope trace shows the voltage directly on the monitor outputs of the heater power supply. The steps are from the voltage resolution of the 4116 DAC.

We also measured the voltage noise on the monitor plugs on the front panel. If these are a true representation of the voltage noise which supplies the heater jacket, then we can use it to estimate the temperature fluctuations of the can. Using the spectrum of temperature fluctuations, we can estimate the actual length changes of the reference cavity.

I used the new fax/scanner/toaster that Steve and Bob both love to scan this HP spectrum analyzer image directly to a USB stick! It can automatically make PDF from a piece of paper.

The pink trace is the analyzer noise with a 50 Ohm term. The blue trace is the heater supply with the servo turned off. With the servo on (as in the scope trace above) the noise is much much larger because of the DAC steps.

Attachment 1: 09080901.PDF
09080901.PDF
  1977   Tue Sep 8 19:36:52 2009 JenneOmnistructureDMFDMF restarted

I (think I) restarted DMF.  It's on Mafalda, running in matlab (not the complied version which Rana was having trouble with back in the day).  To start Matlab, I did "nohup matlab", ran mdv_config, then started seisBLRMS.m running.  Since I used nohup, I then closed the terminal window, and am crossing my fingers in hopes that it continues to work.  I would have used Screen, but that doesn't seem to work on Mafalda.

  1976   Tue Sep 8 19:30:33 2009 ranaUpdatePSLc1psl rebooted for new RCPID database settings

The RC thermal PID is now controllable from its own MEDM screen which is reachable from the FSS screen. The slowpid.db and psl.db have been modified to add these records and all seems to be working fine.

Also, I've attached the c1psl startup output that we got on the terminal. This is just for posterity.

I'm also done tuning the PID for now. Using Kp = -1.0, Ki = -0.01, and Kd = 0, the can servo now has a time constant of ~10 minutes and good damping as can be seen in the StripTool snap below. These values are also now in the saverestore.req so hopefully its fully commissioned.

I bet that its much better now than the MINCO at holding against the 24 hour cycle and can nicely handle impulses (like when Steve scans the table). Lets revisit this in a week to see if it requires more tuning.

Attachment 1: c1psl-term-dump.txt.gz
Attachment 2: C1PSL_FSS_RCPID.png
C1PSL_FSS_RCPID.png
Attachment 3: Picture_1.png
Picture_1.png
  1975   Tue Sep 8 17:57:30 2009 JenneUpdatePEMAll the Acc/Seis working again

All of the accelerometers and seismometers are plugged in and functional again.  The cables to the back of the accelerometer preamp board (sitting under the BS oplev table) had been unplugged, which was unexpected.  I finally figured out that that's what the problem was with half of the accelerometers, plugged them back in, and now all of the sensors are up and running.

TheSEIS_GUR seismometer is under MC1, and all the others (the other Guralp, the Ranger which is oriented vertically, and all 6 accelerometers) are under MC2.

  1974   Tue Sep 8 16:01:52 2009 steveConfigurationVACRGA was separeted from IFO

The RGA isolation valve VM1 was closed since Aug 24, 2009  I installed the new UPS that time.

The last RGA scan in the log is from Aug 7, 2009   The vacuum rack UPS failed on Aug 15, 2009

I opened VM1 today so we can have ifo rga scan tomorrow.

 

Attachment 1: vm1closed.jpg
vm1closed.jpg
  1973   Tue Sep 8 15:14:26 2009 rana, alexConfigurationDAQRAID update to Framebuilder: directories added + lookback increased

 Alex logged in around 10:30 this morning and, at our request, adjusted the configuration of fb40m to have 20 days of lookback.

I wasn't able to get him to elog, but he did email the procedure to us:


1) create a bunch of new "Data???" directories in /frames/full
2) change the setting in /usr/controls/daqdrc file
       set num_dirs=480;

my guess is that the next step is:

3) telnet fb0 8087

    daqd>  shutdown

I checked and we do, in fact, now have 480 directories in /frames/full and are so far using up 11% of our 13TB capacity. Lets try to remember to check up on this so that it doesn't get overfull and crash the framebuilder.

  1972   Tue Sep 8 12:26:16 2009 AlbertoUpdatePSLConnection of the RC heater's power supply replaced

I have replaced the temporary clamps that were connecting the RC heater to its power supply with a new permanent connection.

In the IY1 rack, I connected the control signal of the RC PID temperature servo - C1:PSL-FSS_TIDALSET - to the input of the RC heater's power supply.

The signal comes from a DAC in the same rack, through a pair of wires connected to the J9-4116*3-P3 cross-connector (FLKM). I joined the pair to the wires of the BNC cable coming from the power supply, by twisting and screwing them into two available clamps of the breakout FKLM in the IY1 rack - the same connected to the ribbon cable from RC Tmeperature box.

Instead of opening the BNC cable coming from the power supply, I thought it was a cleaner and more robust solution to use a BNC-to-crocodile clamp from which I had cut the clamps off.

During the transition process, I connected the power supply BNC input to a a voltage source that I set at the same voltage of the control signal before I disconnected it (~1.145V).

I monitored the temperature signals and it looked like the RC Temperature wasn't significantly affected by the operation.

  1971   Mon Sep 7 23:51:48 2009 ranaConfigurationComputersmatlab installed: 64-bit linux

I have wiped out the 2008a install of 64-bit linux matlab and installed 2009a in its place. Enjoy.

  1970   Mon Sep 7 23:35:03 2009 ranaUpdatePSLRC thermal servo: PID script modified, database + screen added

I have added the records for the RC thermal PID servo into the psl/slowpid.db file which also holds the records for the SLOW servo that uses the NPRO-SLOW to minimize the NPRO-FAST. This new database will take effect upon the next PSL boot.

The perl script which runs the servo is scripts/PSL/FSS/RCthermalPID.pl. Right now it is using hard-coded PID parameters - I will modify it to use the on-screen values after we reboot c1psl.

The new screen C1PSL_FSS_RCPID.adl, the script, and the .db have been added to the SVN.

I have got some preliminary PID parameters which seem to be pretty good: The RCTEMP recovers in ~10 minutes from a 1 deg temperature step and the closed loop system is underdamped with a Q of ~1-2.

I'm leaving it running on op340m for now - if it goes crazy feel free to do a 'pkill RCthermalPID.pl'.

Attachment 1: Untitled.png
Untitled.png
  1969   Mon Sep 7 23:18:01 2009 AlbertoUpdateLockingSome locking attempts

Tried to lock the interferometer but arm power didn't get over 65.

Tonight, after the weekend, I resumed the work on locking.

When I started the Mode Cleaner was unlocked because the MZ was also unlocked.

I aligned the MZ and the transmitted power reached about 2.5

Initially the interferometer lost lock at arm power of about 3-4. It looked like the alignment wasn't good enough. So I ran the alignment scripts a few times, first the scripts for the single parts and in the end the one for the full IFO.

Then I also locked again the MZ and this time the transmitted power got to about 4.

In the following locking attempts the the arm power reached 65 but then the lock got lost during the handing of CARM to C1:LSC-PD11_I

I'll keep working on that tomorrow night.

  1968   Mon Sep 7 20:05:18 2009 ranaUpdatePSLRCTEMP v. RMTEMP

Since ~Aug. 27, the reference cavity has been running with no thermal control. This is not really a problem at the 40m; a 1 deg change of the glass cavity

will produce a 5 x 10-7 strain in the arm cavity. That's around 20 microns of length change.

This open loop time gave us the opportunity to see how good our cavity's vacuum can insulation is.

 

rct.pngrct2.png

The first plot below shows the RCTEMP sensors and the RMTEMP sensor. RMTEMP is screwed down to the table close to the can and RCTEMP is on the can, underneath the insulation. I have added a 15 deg offset to RMTEMP so that it would line up with RCTEMP and allow us to see, by eye, what's happening.

There's not enough data here to get a good TF estimate, but if we treat the room temperature as a single frequency (1 / 24 hours) sine wave source, then we can measure the delay and treat it as a phase shift. There's a ~3 hour delay between the RMTEMP and RCTEMP. If the foam acts like a single pole low pass filter, then the phase delay of (3/24)*360 = 45 deg implies a pole at a ~3 hour period. I am not so sure that this is a good foam model, however.

The colorful plot is a scatter plot of RCTEMP v. RMTEMP. The color denotes the time axis - it starts out blue and then becomes red after ten days.

  1967   Fri Sep 4 16:09:26 2009 josephbSummaryVACRebooted RGA computer and reset RGA settings

Steve noticed the RGA was not working today.  It was powered on but no other lights were lit.

Turns out the c0rga machine had not been rebooted when the file system on linux1 was moved to the raid array, and thus no longer had a valid mount to /cvs/cds/.  Thus, the scripts that were run as a cron could not be called.

We rebooted c0rga, and then ran ./RGAset.py to reset all the RGA settings, which had been reset when the RGA had lost power (and thus was the reason for only the power light being lit).

 

Everything seems to be working now.  I'll be adding c0rga to the list of computers to reboot in the wiki.

  1966   Thu Sep 3 23:41:32 2009 AlbertoConfigurationLSCPOX (PD3) aligned

Today I aligned the beam to PD3 (POX) since Steve had moved it.

The DC power read 1.3mV when the beam was on the PD.

  1965   Thu Sep 3 11:20:26 2009 steveUpdateGeneralelectrical ground in place

Quote:

I was told yesterday, that on Friday the construction people accidentally ripped out one of the 40m soil ground.....AND HOW MANY MORE ARE THERE? nobody knows.

It was ~8 ft long and 0.5" diameter buried in the ground. There is no drawing found to identify this exact building ground. They promised to replace this on Wednesday with a 10 ft long and 0.75" diameter.

The  the wall will be resealed where the conduit enters  the  north west corner of the IFO room 104

 

There should be no concern about safety because the 40m building main ground is connected to the CES Mezzanine step-down transformer.

 Atm1 is showing ground bus under N-breaker panel in 40m IFO room north-west corner.

The second ground bus is visible farther down south under M-breaker panel.

Atm2 is the new ground that will be connected to ground bus-N

Attachment 1: Gs-n.JPG
Gs-n.JPG
Attachment 2: newground.JPG
newground.JPG
  1964   Thu Sep 3 10:19:35 2009 steve, albertoUpdatePEMparticle counts and burning hillsides

The San Gabriel mountain has been on fire for  6 days. 144,000 acres of beautiful hillsides burned down and it's still burning.  Where the fires are.

The 40m lab particle counts are more effected by next door building-gardening activity than the fire itself.

This 100 days plot shows that.

Attachment 1: fire6d.jpg
fire6d.jpg
Attachment 2: 40mgarden.JPG
40mgarden.JPG
Attachment 3: 40mgarden2.JPG
40mgarden2.JPG
Attachment 4: RBLOG-FIRE-SKY.JPG
RBLOG-FIRE-SKY.JPG
  1963   Tue Sep 1 13:52:06 2009 steveConfigurationGeneralPOX rfpd is back & needs alignment

Quote:

I removed POX rfpd to see how it is mounted on its base. It is here on the work bench just in case someone wants to use it the IFO over the week end.

 I put POX back to it's place with markers. The pd was removed from it's base so it is for sure misaligned.

  1962   Tue Sep 1 11:23:36 2009 steveUpdateGeneralelectrical ground

I was told yesterday, that on Friday the construction people accidentally ripped out one of the 40m soil ground.....AND HOW MANY MORE ARE THERE? nobody knows.

It was ~8 ft long and 0.5" diameter buried in the ground. There is no drawing found to identify this exact building ground. They promised to replace this on Wednesday with a 10 ft long and 0.75" diameter.

The  the wall will be resealed where the conduit enters  the  north west corner of the IFO room 104

 

There should be no concern about safety because the 40m building main ground is connected to the CES Mezzanine step-down transformer.

Attachment 1: ground.JPG
ground.JPG
  1961   Fri Aug 28 15:30:15 2009 steveConfigurationGeneralPOX rfpd removed

I removed POX rfpd to see how it is mounted on its base. It is here on the work bench just in case someone wants to use it the IFO over the week end.

  1960   Fri Aug 28 13:49:07 2009 robUpdateLockingRF CARM hand off problem

Quote:
Last night, the lock script proceeded to the RF CARM hand-off about half of the time.
However, the hand off was still unsuccessful.

It failed instantly when you turn on the REFL1 input of the CM board, even
when the REFL1 input gain was very low, like -28dB.

I went to the LSC rack and checked the cabling.
The output from the PD11_I (REFL_2) demodulation board is split
into two paths. One goes directly to the ADC and the other one goes
to an SR560. This SR560 is used just as an inverter. Then
the signal goes to the REFL1 input of the CM board.

I found that the SR560 was set to the A-B mode, but B input was open.
This made the signal very noisy. So I changed it to A only mode.
There was also a 1/4 attenuator between the PD11_I output and the SR560.
I took it out and reduced the gain of SR560 from 10 to 2.
These changes allowed me to increase the REFL1 gain to -22dB or so.
But it is still not enough.

I wanted to check the CM open loop TF before the hand-off, but I could
not do that because the lock was lost instantly as soon as I enabled the
test input B of the CM board.
Something is wrong with the board ?

Using the PD11_I signal going into the ADC, I measured the transfer functions
from the CM excitation (digital one) to the REFL_DC (DC CARM signal) and PD11_I.
The TF shapes matched. So the PD11_I signal itself should be fine.

We should try:
* See if flipping the sign of PD11_I signal going to REFL1 input solve the problem.
* Try to measure the CM analog TF again.
* If the noise from the servo analyzer is a problem, try to increase the input gains
of the CM board and reduce the output gain accordingly, so that the signal flowing
inside the CM board is larger.



I'd bet it's in a really twitchy state by the time the script gets to the RF CARM handoff, as the script is not really validated up to that point. It's just the old script with a few haphazard mods, so it needs to be adjusted to accomodate the 15% power drop we've experienced since the last time it was locked.

The CM servo gain needs to be tweaked earlier in the script--you should be able to measure the AO path TF with the arm powers at 30 or so. I was able to do this with the current SR785 setup earlier this week without any trouble.

The 1/4 attenuator is there to prevent saturations on the input to the SR560 when there's still a CARM offset.

Not sure if flipping the sign of PD11 is right, but it's possible we compensated the digital gains and forgot about it. This signal is used for SRCL in the initial acquisition, so we'd have noticed a sign flip.
  1959   Fri Aug 28 12:56:17 2009 YoichiUpdateLockingRF CARM hand off problem
Last night, the lock script proceeded to the RF CARM hand-off about half of the time.
However, the hand off was still unsuccessful.

It failed instantly when you turn on the REFL1 input of the CM board, even
when the REFL1 input gain was very low, like -28dB.

I went to the LSC rack and checked the cabling.
The output from the PD11_I (REFL_2) demodulation board is split
into two paths. One goes directly to the ADC and the other one goes
to an SR560. This SR560 is used just as an inverter. Then
the signal goes to the REFL1 input of the CM board.

I found that the SR560 was set to the A-B mode, but B input was open.
This made the signal very noisy. So I changed it to A only mode.
There was also a 1/4 attenuator between the PD11_I output and the SR560.
I took it out and reduced the gain of SR560 from 10 to 2.
These changes allowed me to increase the REFL1 gain to -22dB or so.
But it is still not enough.

I wanted to check the CM open loop TF before the hand-off, but I could
not do that because the lock was lost instantly as soon as I enabled the
test input B of the CM board.
Something is wrong with the board ?

Using the PD11_I signal going into the ADC, I measured the transfer functions
from the CM excitation (digital one) to the REFL_DC (DC CARM signal) and PD11_I.
The TF shapes matched. So the PD11_I signal itself should be fine.

We should try:
* See if flipping the sign of PD11_I signal going to REFL1 input solve the problem.
* Try to measure the CM analog TF again.
* If the noise from the servo analyzer is a problem, try to increase the input gains
of the CM board and reduce the output gain accordingly, so that the signal flowing
inside the CM board is larger.
  1958   Thu Aug 27 16:14:28 2009 steveSummaryOMCburned photodiode

Old -pre 6/2009  LLO DCPD 3 mm od GTRAN photodiode

Attachment 1: 20090827_173252.jpg
20090827_173252.jpg
Attachment 2: 20090827_170802.jpg
20090827_170802.jpg
  1957   Thu Aug 27 14:00:33 2009 ranaUpdatePSLRC thermal servo impulse response

I stepped the TIDALSET and looked at what happened. Loop was closed with the very low gain.

The RED guy tells us the step/impulse response of the RC can to a step in the heater voltage.

The GREY SLOWDC tells us how much the actual glass spacer of the reference cavity lags the outside can temperature.

Since MINCOMEAS is our error signal, I have upped his SCAN period from 0.5 to 0.1 seconds in the database and reduced its SMOO from 0.9 to 0.0. I've also copied over the Fricke SLOW code and started making a perl PID loop for the reference cavity.

Attachment 1: Untitled.png
Untitled.png
  1956   Thu Aug 27 13:42:08 2009 ranaSummaryPSLReference Cavity Temperature Control: psl.db changes

I made the changes to the psl.db to handle the new Temperature box hardware. The calibrations (EGUF/EGUL) are just copied directly from the LHO .db file (I have rsync'd their entire target area to here).

allegra:c1psl>diff psl.db~ psl.db
341,353d340
< grecord(ai,"C1:PSL-FSS_TIDALOUT")
< {
<       field(DESC,"TIDALOUT- drive to the reference cavity heater")
<       field(DISV,"1")
<         field(SCAN,".5 second")
<       field(DTYP,"VMIVME-3113")
<       field(INP,"#C0 S28 @")
<       field(EGUF,"10")
<       field(EGUL,"-10")
<       field(EGU,"volts")
<       field(LOPR,"-10")
<       field(AOFF,"0")
< }
493,494c480,481
<         field(EGUF,"285.675")
<         field(EGUL,"-214.325")
---
>         field(EGUF,"67.02")
>         field(EGUL,"7.96")
508,509c495,496
<         field(EGUF,"726.85")
<         field(EGUL,"-1273.15")
---
>         field(EGUF,"75.57")
>         field(EGUL,"12.31")
531,532c518,519
<         field(EGUF,"726.85")
<         field(EGUL,"-1273.15")
---
>         field(EGUF,"75.57")
>         field(EGUL,"12.31")
605,617d591
< grecord(ai,"C1:PSL-FSS_TIDALINPUT")
< {
<       field(DESC,"TIDALINPUT- tidal actuator input")
<       field(DISV,"1")
<         field(SCAN,".5 second")
<       field(DTYP,"VMIVME-3123")
<       field(INP,"#C0 S3 @")
<       field(EGUF,"10")
<       field(EGUL,"-10")
<       field(EGU,"volts")
<       field(LOPR,"-10")
<       field(AOFF,"0")
< }
1130a1105,1130
> grecord(ai,"C1:PSL-FSS_TIDALINPUT")
> {
>       field(DESC,"TIDALINPUT- tidal actuator input")
>       field(DISV,"1")
>         field(SCAN,".5 second")
>       field(DTYP,"VMIVME-3123")
>       field(INP,"#C0 S3 @")
>       field(EGUF,"10")
>       field(EGUL,"-10")
>       field(EGU,"volts")
>       field(LOPR,"-10")
>       field(AOFF,"0")
> }
> grecord(ai,"C1:PSL-FSS_TIDALOUT")
> {
>       field(DESC,"TIDALOUT- drive to the reference cavity heater")
>       field(DISV,"1")
>         field(SCAN,".5 second")
>       field(DTYP,"VMIVME-3113")
>       field(INP,"#C0 S28 @")
>       field(EGUF,"10")
>       field(EGUL,"-10")
>       field(EGU,"volts")
>       field(LOPR,"-10")
>       field(AOFF,"0")
> }
1143,1144c1143,1144
<         field(HOPR,"0.010")
<         field(LOPR,"-0.010")
---
>         field(HOPR,"2")
>         field(LOPR,"0")

  1955   Thu Aug 27 12:34:48 2009 YoichiUpdateLockingup to arm power 70
Last night, I tried to run locking scripts.
The power went up to 70 a couple of times .
Then it failed to switch to RF CARM.
I was too tired at that time to figure out what is the problem with the switching.
But it seemed to me that the problem could be solved by some gain tweaking.
Looks like the IFO is back to a good state.
  1954   Wed Aug 26 19:58:14 2009 Rana, AlbertoUpdatePSLReference Cavity Temperature Control: MINCO PID removed

Summary: This afternoon we managed to get the temperature control of the reference cavity working again.

We bypassed the MINCO PID by connecting the temperature box error signal directly into EPICS.

We couldn't configure the PID so that it worked with the modified temperature box so we decided to just avoid using it.

Now the temperature control is done by a software servo by using the channel C1:PSL-FSS_MINCOMEAS as error signal and driving C1:PSL-FSS_TIDALSET (which we have clip-doodle wired directly to the heater input).

 

We 'successfully' used ezcaservo to stabilize the temperature:

ezcaservo -r C1:PSL-FSS_MINCOMEAS -s 26.6 -g -0.00003 C1:PSL-FSS_TIDALSET

 

We also recalibrated the channels:

C1:PSL-FSS_RMTEMP

C1:PSL-FSS_RCTEMP

C1:PSL-FSS_MINCOMEAS

with Peter King on the phone by using ezcawrite (EGUF and EGUL) but we didn't change the database yet. So please do not reboot the PSL computer until we update the database.

 

More details will follow.

Attachment 1: rc.png
rc.png
  1953   Wed Aug 26 16:35:03 2009 AlbertoConfigurationPSLPSL reference cavity temperature box modifications

Basically, in addition to the replacement of the resistors with metal film ones, Peter replaced the chip that provides a voltage reference.

The old one provided about 2.5 V, whereas the new one gets to about 7V. Such reference voltage somehow depends on the room temperature and it is used to generate an error signal for the temperature of the reference cavity.

Peter said that the new higher reference should work better.

  1952   Wed Aug 26 16:31:34 2009 steveUpdatePSLreference cavity temp box temporarly out of order

Quote:

It turned out that half an hour was too long. In less than that the reference cavity temperature passed the critical point when the temperature controller (located just below the ref cav power supply in the same rack) disables the input power to the reference cavity power supply.

The controller's display in the front shows two numbers. The first goes with the temperature of the reference cavity; the second is a threshold set for the first number. The power supply gets enabled only when the first number comes under the threshold value.

Now the cavity is cooling down and it will take about another hour for its temperature to be low enough and for the heater power supply to be powered.

 The cavity temp cooled below SP2 set point 0.1  The Minco SP1 (present temp in Volts) now reading -0.037 so DC power supply was turned on and set to 12V 1A  

 

 

ELOG V3.1.3-